# WhiSPA: Semantically and Psychologically Aligned Whisper with Self-Supervised Contrastive and Student-Teacher Learning

### v2

@@ 1 Introduction
Figure 1: Speech processing pipelines that use text-based representations from language models often yield higher accuracies than those produced solely by SotA audio models. While both approaches ingest the same raw audio signal, we close the performance gap by introducing a speech encoder with similar performance to a text-based LM pipeline.
Human communication is inherently multimodal, but AI integration of modalities is often fragmented Lazaro et al. (2021); Gu et al. (2017), where speech models, such as Whisper Radford et al. (2022), are often pipelined into text-based language models (LMs) Chuang et al. (2020) in order to get the most accurate speech-based representations (see Figure 1). Text-based LMs produce richer semantic representations Wu et al. (2024); Fu (2024). This often results in redundant computational costs from having two LMs in the pipeline (one within the audio model and one for the text LM) and representations remain incomplete of the full spectrum of human expressions Zhang et al. (2023); Lian et al. (2023). y important for psychological and social scientific applications of LMs where reptions from text-based LMs demonstrate superior performance than direct speech repdirect speechchen2024textbetteLukac2024-tb; rspeech.
@@
Our focus on psychological or human-level tasks reflects a growing demand for foundation models to better understand the intrinsic qualities of human communication Soni et al. (2024). As  Clark and Schober (1992) put it, “The common misconception is that language has to do with words and what they mean. It does not. It has to do with people and what they mean.” and specifically how well the representations capture information about the people communicating Hovy and Yang (2021); Soni et al. (2022). More specifically, psychological studies have suggested mental health attributes are highly multimodal as they are influenced by subtle nuances in voice and content Sartori and Orrù ; Chen et al. (2024).
@@ 2 Background
This work builds on top of Whisper Radford et al. (2022), OpenAI’s SotA automatic speech recognition (ASR) foundation model. We chose Whisper over other alternatives such as HuBERT and Wav2Vec2-BERT, since previous works Kyung et al. (2024); Yang et al. (2023) have shown that Whisper has a stronger language encoding module at capturing speaker attributes.
@@
Multi-level fusion architectures leveraging both acoustic and lexical features have shown to improve the performance on downstream tasks. For instance, Zhao et al. (2022) demonstrates that coattention-based early fusion and late fusion using Wav2Vec2.0 Baevski et al. (2020); Schneider et al. (2019) and BERT Devlin et al. (2019) outperform SotA emotion recognition benchmarks.Other recent works inject acoustic nuances into language models using textual descriptions of speech characteristics Wu et al. (2024) or common-sense reasoning through historical utterances from the speaker Fu (2024).However, this approach does not fully leverage the cross-modal dependencies between text and audio, as it remains unimodal, relying solely on textual inputs rather than raw acoustic representations.
@@ 3 Data & Tasks
For each audio/text pair in our datasets, we extract theoretically derived psychological features using pre-trained lexica V Ganesan et al. (2022), which we refer to as PsychEmb. PsychEmb broadly covers three domains of psychological constructs measured at different temporal granularity: (a) states, which reflect the emotional condition of the person at a point in time; (b) dispositions, which are slightly more stable than states and reflect the tendencies of humans to behave in certain ways and finally (c) the traits, which are long term stable characteristics Park et al. (2014). The ten dimensions of PsychEmb are Valence (VAL), Arousal (ARO), Openness (OPE), Consciousness (CON), Extraversion (EXT), Agreeableness (AGR), Neuroticism (NEU), Anger (ANG), Anxiety (ANX), and Depression (DEP), each represented with scalar values. Once the self-supervised PsychEmb dimensions were extracted for each segment across both datasets, we perform a 80:10:10 (train/val/test) split.
@@ 4.1.1 Cosine Similarity Loss (CS)
where $\text{where sim}(\mathbf{A}_{i},\mathbf{T}_{i})=\frac{\mathbf{A}_{i}\cdot\mathbf{T}_{i}}{||\mathbf{A}_{i}||\text{ }||\mathbf{T}_{i}||}$
@@ 4.3 Adding Psychological Alignment (WhiSPA)
Dataset Model Traits States Dispositions OPE CON EXT AGR NEU VAL ARO ANG ANX DEP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ HiTOP W2V2B .63 .14 .69 .12 .75 .09 .60 .09 .72 .10 .65 .001 .73 .000 .45 .04 .51 .02 .64 .03 HuBERT .67 .13 .71 .11 .77 .08 .57 .10 .70 .11 .66 .001 .73 .000 .48 .04 .48 .02 .58 .04 Whisper-384 .74 .11 .80 .08 .69 .10 .76 .06 .78 .08 .71 .001 .82 .000 .53 .03 .61 .01 .65 .03 SBERT-384 .73 .11 .83 .07 .68 .11 .75 .06 .77 .09 .69 .001 .81 .000 .59 .03 .60 .01 .61 .04 WhiSA-384 .71* .11 .81* .08 .70 .10 .77* .06 .78* .08 .73* .001 .83* .000 .59$\dagger$ .03 .61 .01 .61 .04 WhiSPA-384r .74* .11 .83$\dagger$ .07 .70 .10 .79$\dagger$ .05 .79$\dagger$ .07 .78$\dagger$ .000 .85$\dagger$ .000 .59$\dagger$ .03 .61$\dagger$ .01 .66* .03 WhiSPA-394 .72* .11 .83$\dagger$ .07 .72 .09 .79$\dagger$ .05 .82$\dagger$ .07 .76$\dagger$ .000 .84* .000 .62$\dagger$ .03 .65$\dagger$ .01 .63* .03 WTC W2V2B .33 .54 .51 .55 .34 .64 .37 .46 .34 .64 .32 .004 .51 .005 .31 .21 .14 .16 .22 .15 HuBERT .35 .54 .57 .50 .39 .61 .44 .43 .42 .60 .38 .003 .53 .005 .36 .20 .15 .16 .22 .16 Whisper-384 .57 .43 .70 .37 .68 .38 .64 32 .67 .40 .56 .003 .82 .002 .54 .16 .46 .13 .45 .13 SBERT-384 .65 .35 .78 .29 .73 .33 .73 .25 .73 .34 .62 .003 .86 .002 .62 .14 .56 .11 .59 .11 WhiSA-384 .70$\dagger$ .31 .82$\dagger$ .24 .75$\dagger$ .32 .76$\dagger$ .23 .77$\dagger$ .30 .67$\dagger$ .002 .85$\dagger$ .002 .66$\dagger$ .13 .61$\dagger$ .10 .61$\dagger$ .10 WhiSPA-384r .71$\dagger$ .29 .82$\dagger$ .24 .74$\dagger$ .30 .76$\dagger$ .20 .76$\dagger$ .27 .68$\dagger$ .002 .85$\dagger$ .002 .67$\dagger$ .01 .61$\dagger$ .09 .61$\dagger$ .09 WhiSPA-394 .72$\dagger$ .28 .83$\dagger$ .22 .76$\dagger$ .29 .79$\dagger$ .19 .79$\dagger$ .26 .70$\dagger$ .002 .86$\dagger$ .002 .69$\dagger$ .11 .64$\dagger$ .09 .66$\dagger$ .09
@@ 5 Results & Discussion
Semantic-Psychological alignment is SotA for audio-based psychological assessments.
@@
The last row in Table 5 underscores the marginal increase in correlations after appending SBERT embeddings to WhiSPA. WhiSPA, trained through a student-teacher alignment paradigm, appears to reach a semantic and psychological optimum during convergence. This is evident in its substantial performance gains over Whisper, which lacks the semantic and psychological depth provided by language models. However, the potential of cross-modal alignment may be constrained by the representational efficacy of the teacher model(s). On human-annotated audio segments, all of the WhiSPA variants achieve substantial improvements in capturing acoustic valence. In comparison with Whisper-384, WhiSPA-384r exhibits a gain of $+15$ Pearson points in VAL (acoustic valence) which exemplifies the reduction in the semantic/psychological gap between audio models and text-based models. Notably in Figure 8, WhiSPA-394 demonstrates clear improvements in specific measures such as INT and VAL, with gains of $+3$ and $+8$ Pearson points, respectively, when compared to its teacher, SBERT-384.
@@
6(a) shows that n-grams known to be indicative of PTSD severity from prior studies Kjell et al. (2024) —- including first-person pronouns, experienced symptoms, psychological distress, and negative affect – yield significantly higher correlations with WhiSPA’s predictions compared to Whisper. In contrast, 6(b) reveals that language discussing relationships and positive affect is more negatively associated with WhiSPA’s scores. These findings indicate that the contrastive loss training effectively aligns the latent space with rich semantic and psychological representations, capturing psychologically relevant linguistic markers more robustly. The highly semantic latent spaces of text-based LMs are reflected in WhiSPA’s representations, especially for psychological nuances in spoken language. More quantitative analysis of our model can be found in Appendix subsection A.4
@@ 8 Ethical Implications
The WTC and HiTOP recordings took place in a clinical setting where each participant gave consent and was fully informed about the study, that it was voluntary to take part, and that they had the right to withdraw at any time without giving a reason or that it would affect their treatment. After the interview, participants were debriefed (for more details about the WTC data collection, see Kjell et al. (2024); for more details about the HiTOP data, see Kotov et al. (2022, 2024). The studies and data uses were approved by the Institutional Review Board at an undisclosed university for privacy reasons.
@@ Acknowledgments
The work presented in this paper stems from the immense hours of audio recordings from the WTC and HiTOP datasets. We greatly thank the participants, creators, maintainers, and interviewers from the studies for enabling this research.
@@ A.1 Data Description
The HiTOP dataset consists of video-recorded interviews conducted between World Trade Centre responder participants and clinicians. Each recording is annotated with the outcomes derived from the HiTOP structured interview, which includes a standardized set of questions designed to assess a comprehensive set of mental health dimensions, including aspects of internalizing (e.g., questions about distress and fear), dis-inhibited externalizing (e.g., questions about substance abuse and antisocial behaviours) and more.
@@ A.2 Training
Additionally, we use open-source licensed pre-trained models from HuggingFace. Our programmatic implementation for deep learning is done with PyTorch. When it comes to evaluation, we utilize Differential Language Analysis Tool Kit (DLATK) for correlating regression results across specified groups (i.e., $user\_id$ or $segment\_id$)
@@
Cosine similarity is sensitive to the relative magnitudes of the vectors being compared. If the added ten dimensions of psychological features have a very different scale or distribution from SBERT embeddings, they could dominate or skew the cosine similarity computation. Once either loss function is applied, (1) or (2), WhiSPA embeddings remain semantically aligned with SBERT while also encoding meaningful affective cues for downstream tasks.
@@
During the training of WhiSPA, we experimented with identifying which dimensions of the teacher-model, SBERT, have the lowest correlations with PsychEmb dimensions to replace those dimensions. We decided that this approach may lead to statistical biases when training, and so we naively replaced the first 10 dimensions. One should note that the set of 10 dimensions to replace in SBERT can be chosen arbitrarily since our study experimented with this.

### v4

@@ 1 Introduction
Human communication is inherently multimodal, but AI integration of modalities is often fragmented Lazaro et al. (2021); Gu et al. (2017). Speech to text models, like Whisper Radford et al. (2022), are often pipelined into text-based language models (LMs) Chuang et al. (2020) in order to get the most accurate speech-based representations (see Figure 1).
Figure 1: Speech processing pipelines that are further processed by language models (option #1 above) often yield higher accuracies than those produced solely by SotA audio models (option #2). Our distillation introduces a speech encoder which streamlines the pipeline and exhibits similar performance to a text-based LM.
This often results in redundant computational costs from having two LMs in the pipeline (one within the audio model and one for the text LM) and representations remain incomplete of the full spectrum of human expressions Zhang et al. (2023); Lian et al. (2023). This is especially important for psychological and social scientific applications where representations from text-based LMs demonstrate superior performance than direct speech representations Lukac (2024); Chen et al. (2024).
@@
Our focus on psychological or human-level tasks reflects a growing demand for foundation models to better understand the intrinsic qualities of human communication Soni et al. (2024). As  Clark and Schober (1992) put it, “The common misconception is that language has to do with words and what they mean. It does not. It has to do with people and what they mean.” and specifically how well the representations capture information about the people communicating Hovy and Yang (2021); Soni et al. (2022). More specifically, psychological studies have suggested mental health attributes are highly multimodal as they are influenced by subtle nuances in voice and content Sartori and Orrù (2023); Chen et al. (2024).
@@ 2 Background
This work builds on top of Whisper Radford et al. (2022), OpenAI’s SotA automatic speech recognition (ASR) foundation model. We chose Whisper over other alternatives such as HuBERT and Wav2Vec2-BERT, since previous works Kyung et al. (2024); Yang et al. (2023) have shown that Whisper has a stronger language encoding module at capturing speaker attributes due to its pretraining objective of transcription/translation.
@@
Multi-level fusion architectures leveraging both acoustic and lexical features have shown to improve performance in emotion recognition, speaker identification, and other downstream tasks. For instance, Zhao et al. (2022) demonstrates that coattention-based early fusion and late fusion using Wav2Vec2.0 Baevski et al. (2020); Schneider et al. (2019) and BERT Devlin et al. (2019) outperform SotA emotion recognition benchmarks. Other recent works inject acoustic nuances into language models using textual descriptions of speech characteristics Wu et al. (2024) or common-sense reasoning through historical utterances from the speaker Fu (2024). However, this approach does not fully leverage the cross-modal dependencies between text and audio, as it remains unimodal, relying solely on textual inputs rather than raw acoustic representations.
@@ 3 Data & Tasks
For each audio/text pair in our datasets, we extract theoretically derived psychological features using pre-trained lexica V Ganesan et al. (2022), which we refer to as PsychEmb. PsychEmb broadly covers three domains of psychological constructs measured at different temporal granularity: (a) states, which reflect the emotional condition of the person at a point in time; (b) dispositions, which are slightly more stable than states and reflect the tendencies of humans to behave in certain ways and finally (c) the traits, which are long term stable characteristics Park et al. (2014). The ten dimensions of PsychEmb are Valence (VAL), Arousal (ARO), Openness (OPE), Consciousness (CON), Extraversion (EXT), Agreeableness (AGR), Neuroticism (NEU), Anger (ANG), Anxiety (ANX), and Depression (DEP), each represented with scalar values. After extracting self-supervised PsychEmb dimensions for each segment across both datasets, we perform a 80:10:10 (train/val/test) split.
@@ 4.1.1 Cosine Similarity Loss (CS)
$\text{sim}(\mathbf{A}_{i},\mathbf{T}_{i})=\frac{\mathbf{A}_{i}\cdot\mathbf{T}_{i}}{||\mathbf{A}_{i}||\text{ }||\mathbf{T}_{i}||}$
@@ 4.3 Adding Psychological Alignment (WhiSPA)
Dataset Model Traits States Dispositions OPE CON EXT AGR NEU VAL ARO ANG ANX DEP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ SBERT-384 .73 .11 .83 .07 .68 .11 .75 .06 .77 .09 .69 .001 .81 .000 .59 .03 .60 .01 .61 .04 HiTOP W2V2B .63 .14 .69 .12 .75 .09 .60 .09 .72 .10 .65 .001 .73 .000 .45 .04 .51 .02 .64 .03 HuBERT .67 .13 .71 .11 .77 .08 .57 .10 .70 .11 .66 .001 .73 .000 .48 .04 .48 .02 .58 .04 Whisper-384 .74 .11 .80 .08 .69 .10 .76 .06 .78 .08 .71 .001 .82 .000 .53 .03 .61 .01 .65 .03 WhiSA-384 .71* .11 .81* .08 .70 .10 .77* .06 .78* .08 .73* .001 .83* .000 .59$\dagger$ .03 .61 .01 .61 .04 WhiSPA-384r .74* .11 .83$\dagger$ .07 .70 .10 .79$\dagger$ .05 .79$\dagger$ .07 .78$\dagger$ .000 .85$\dagger$ .000 .59$\dagger$ .03 .61$\dagger$ .01 .66* .03 WhiSPA-394 .72* .11 .83$\dagger$ .07 .72 .09 .79$\dagger$ .05 .82$\dagger$ .07 .76$\dagger$ .000 .84* .000 .62$\dagger$ .03 .65$\dagger$ .01 .63* .03 SBERT-384 .65 .35 .78 .29 .73 .33 .73 .25 .73 .34 .62 .003 .86 .002 .62 .14 .56 .11 .59 .11 WTC W2V2B .33 .54 .51 .55 .34 .64 .37 .46 .34 .64 .32 .004 .51 .005 .31 .21 .14 .16 .22 .15 HuBERT .35 .54 .57 .50 .39 .61 .44 .43 .42 .60 .38 .003 .53 .005 .36 .20 .15 .16 .22 .16 Whisper-384 .57 .43 .70 .37 .68 .38 .64 32 .67 .40 .56 .003 .82 .002 .54 .16 .46 .13 .45 .13 WhiSA-384 .70$\dagger$ .31 .82$\dagger$ .24 .75$\dagger$ .32 .76$\dagger$ .23 .77$\dagger$ .30 .67$\dagger$ .002 .85$\dagger$ .002 .66$\dagger$ .13 .61$\dagger$ .10 .61$\dagger$ .10 WhiSPA-384r .71$\dagger$ .29 .82$\dagger$ .24 .74$\dagger$ .30 .76$\dagger$ .20 .76$\dagger$ .27 .68$\dagger$ .002 .85$\dagger$ .002 .67$\dagger$ .01 .61$\dagger$ .09 .61$\dagger$ .09 WhiSPA-394 .72$\dagger$ .28 .83$\dagger$ .22 .76$\dagger$ .29 .79$\dagger$ .19 .79$\dagger$ .26 .70$\dagger$ .002 .86$\dagger$ .002 .69$\dagger$ .11 .64$\dagger$ .09 .66$\dagger$ .09
@@ 5 Results & Discussion
Semantic-Psychological alignment is SotA for speech-based psychological assessments.
@@
Model HiTOP WTC INT DIS ANT SOM THD DET PCL REX AVO NAM HYP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ Transcribe $\rightarrow$ Text Model (Upperbound) SBERT-384 .54* .16 .55* .19 .43* .10 .16* .23 .40* .10 .40* .18 .36* 118.21 .32 11.08 .24* 3.80 .32* 10.08 .33* 15.43 SBERT-1024 .65* .13 .59* .17 .51* .09 .20 .23 .43* .09 .44* .18 .37* 118.67 .32* 10.99 .27* 3.68 .31* 10.26 .31* 15.66 Audio Model Whisper-384 .39 .19 .33 .24 .33 .11 .07 .23 .28 .11 .29 .20 .23 128.85 .21 11.77 .06 4.00 .19 10.87 .23 16.41 WhiSPA-394 .57* .15 .54* .19 .43* .10 .22* .22 .37* .10 .38* .19 .35* 118.91 .30* 11.18 .20 3.85 .32* 10.09 .32* 15.48 Audio Model (Scaled Up) Whisper-1024 .57 .15 .52 .19 .44 .10 .23 .22 .34 .10 .37 .19 .28 126.03 .23 11.81 .11 3.96 .27 10.54 .26 16.29 WhiSPA-1034 .67* .13 .62* .16 .53* .10 .21 .22 .40* .10 .44* .18 .41* 114.44 .34* 10.89 .27* 3.70 .37* 9.84 .34* 15.42
[…]
@@
WhiSPA effectively scales to larger dimensionalities.
To investigate the effects of utilizing a larger teacher LM, we conducted experiments with all-roberta-large-v1 ($\sim$330M) paired with whisper-medium ($\sim$796M) as the audio backbone, each with 1024 embedding dimensions. Table 5 shows that the distillation process remains effective for aligning larger student-teacher model configurations, further validating its scalability and generalizability for the downstream task. When comparing Whisper-384 to WhiSPA-394, we observe an average error reduction of 83.38%, while the 1024-sized models show an even greater reduction of 86.61%. A larger audio backbone also improves the consistency with which the student model outperforms its language-based teacher, likely due to enhanced context retention afforded by a larger parameter space. For example, WhiSPA-394 surpasses its teacher in only 2 of 11 outcomes, whereas WhiSPA-1034 does so in 7 of 11 psychological assessments. These findings underscore the effectiveness of our distillation strategy, particularly for larger models offering greater embedding dimensionality.
@@
The last row in Table 6 underscores the marginal increase in correlations after appending SBERT embeddings to WhiSPA. WhiSPA, trained through a student-teacher alignment paradigm, appears to reach a semantic and psychological optimum during convergence. This is evident in its substantial performance gains over Whisper, which lacks the semantic and psychological depth provided by language models. However, the potential of cross-modal alignment may be constrained by the representational efficacy of the teacher model(s). On human-annotated audio segments, all of the WhiSPA variants achieve substantial improvements in capturing acoustic valence. In comparison with Whisper-384, WhiSPA-384r exhibits a gain of $+15$ Pearson points in VAL (acoustic valence) which exemplifies the reduction in the semantic/psychological gap between audio models and text-based models. Notably in Figure 4, WhiSPA-1034 demonstrates clear improvements in the majority of outcomes, with an average gain of $+2$ Pearson points, when compared to its teacher, SBERT-1024.
@@
Figure 4: WhiSPA Bridges the Semantic/Psychological Representation Gap. WhiSPA consistently outperforms every baseline audio model and, in most cases, matches or exceeds the performance of the text-based language model teacher.
@@
7(a) shows that n-grams known to be indicative of PTSD severity from prior studies Kjell et al. (2024) —- including first-person pronouns, experienced symptoms, psychological distress, and negative affect – yield significantly higher correlations with WhiSPA’s predictions compared to Whisper. In contrast, 7(b) reveals that language discussing relationships and positive affect is more negatively associated with WhiSPA’s scores. These findings indicate that the contrastive loss training effectively aligns the latent space with rich semantic and psychological representations, capturing psychologically relevant linguistic markers more robustly. The highly semantic latent spaces of text-based LMs are reflected in WhiSPA’s representations, especially for psychological nuances in spoken language. More quantitative analysis of our model can be found in Appendix subsection A.4
@@ 8 Ethical Implications
This work was part of a study approved by SBU’s Institutional Review Board (IRB #1157153 World Trade Center Responder Language and Health Study) and (IRB #2022-00391: iHiTOP). The WTC and HiTOP recordings took place in a clinical setting at the Stony Brook WTC Health and Wellness Program where each participant gave consent and was fully informed about the study, that it was voluntary to take part, and that they had the right to withdraw at any time without giving a reason or that it would affect their treatment. After the interview, participants were debriefed (for more details about the WTC data collection, see Kjell et al. (2024); for more details about the HiTOP data, see Kotov et al. (2022, 2024). The studies and data uses were approved by the Institutional Review Board at an undisclosed university for privacy reasons.
@@ Acknowledgments
This work was supported in part by a grant from the CDC/NIOSH (U01 OH012476) and a grant from the NIH-NIAAA (R01 AA028032). The conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied of the CDC, NIH, or any other government organization. We greatly thank the participants, creators, maintainers, and interviewers from the WTC and HiTOP studies for enabling this research.
@@ A.1 Data Description
The HiTOP dataset consists of video-recorded interviews conducted between World Trade Center responder participants and clinicians. Each recording is annotated with the outcomes derived from the HiTOP structured interview, which includes a standardized set of questions designed to assess a comprehensive set of mental health dimensions, including aspects of internalizing (e.g., questions about distress and fear), dis-inhibited externalizing (e.g., questions about substance abuse and antisocial behaviors) and more. This dataset comprises of 525 unique participants.
@@ A.2 Training
Additionally, we use open-source licensed pre-trained models from HuggingFace. Our programmatic implementation for deep learning is done with PyTorch. When it comes to evaluation, we utilize Differential Language Analysis Tool Kit (DLATK) for correlating regression results across specified groups (i.e., $user\_id$ or $segment\_id$).
@@
Cosine similarity is sensitive to the relative magnitudes of the vectors being compared. If the added ten dimensions of psychological features have a very different scale or distribution from SBERT embeddings as visualized in Figure 6, they could dominate or skew the cosine similarity computation. Once either loss function is applied, (1) or (2), WhiSPA embeddings remain semantically aligned with SBERT while also encoding meaningful affective cues for downstream tasks.
@@
During the training of WhiSPA, we experimented with identifying which dimensions of the teacher-model, SBERT, have the lowest correlations with PsychEmb dimensions to replace those dimensions which is depicted in Figure 7. We decided that this approach may lead to statistical biases when training, and so we naively replaced the first 10 dimensions. One should note that the set of 10 dimensions to replace in SBERT can be chosen arbitrarily since our study experimented with this.